OpenStudy (anonymous):

Let it be the matrix A, with the property:

OpenStudy (anonymous):

matrix A: \[\left(\begin{matrix}a & b \\ c & d\end{matrix}\right)\]

OpenStudy (anonymous):

its property: \[A^{2} = O_{2}\]

OpenStudy (anonymous):

solve the equation: \[X^{2} = A\] for a = 1

OpenStudy (anonymous):

\[A^{2} = \left(\begin{matrix}1+bc & b(d+1) \\ c(d+1) & bc+d^{2}\end{matrix}\right)\]

OpenStudy (anonymous):

after solving the equation system i get : \[A = \left(\begin{matrix}1 & \frac{ -1 }{ c } \\ c & -1\end{matrix}\right)\]

OpenStudy (anonymous):

the Trace is 0, same as the determinant.

OpenStudy (turingtest):

what is \(O_2\) ?

OpenStudy (anonymous):

\[\left(\begin{matrix}0 & 0 \\ 0 & 0\end{matrix}\right)\]

OpenStudy (anonymous):

so what'st the question?

OpenStudy (anonymous):

i have to solve the equation : \[X^{2} = \left(\begin{matrix}a & b \\ c & d\end{matrix}\right) \] for a = 1